# Constrained Trajectory Optimization for Hybrid Dynamical Systems

### v1

@@ Abstract
{IEEEkeywords}
Hybrid systems, Constrained control, Optimization
1 Introduction
@@ 2 Related Work
Direct methods transcribe the Optimal Control Problem (OCP) into a Nonlinear Programming (NLP) problem, and have been widely employed in robotic trajectory planning and control. However, hybrid dynamics involving continuous dynamics and discrete events often necessitate Linear Complementarity Programs (LCP) or Mixed Integer Programs (MIP) [9, 13]. Pardo et al. [14] presented an algorithm for legged robots’ locomotion, employing direct collocation within the constraint-consistent subspace defined by contact configurations. This method, incorporating projected impact dynamics constraints, was validated on a hydraulically-actuated quadruped robot. Posa et al. [6] introduced DIRCON, extending direct collocation with manifold constraints and third-order integration accuracy, demonstrating its efficacy in complex locomotion tasks. While these methods predefine a mode or gait sequence, limiting scalability, contact-implicit frameworks have improved optimization capabilities. Patel et al. [13] improved trajectory accuracy by combining direct collocation using higher-order orthogonal polynomials with contact-implicit optimization, eliminating the need for mode scheduling and a priori contact knowledge. Posa et al. [9] addressed trajectory planning with inelastic contact models formulated as LCP, solved via Sequential Quadratic Programming, achieving accurate contact-implicit trajectory optimization. Additionally, Manchester et al. [15] developed algorithms based on discrete variational mechanics for higher-order time-stepping methods, enhancing the accuracy of contact-implicit trajectory optimization.
@@ 4 Hybrid iLQR
$\begin{split}Q_{x}&=\ell_{x}+f_{x}^{\top}\Xi^{\top}V_{x}^{\prime}\\ Q_{u}&=\ell_{u}+f_{u}^{\top}\Xi^{\top}V_{x}^{\prime}\\ Q_{xx}&=\ell_{xx}+f_{x}^{\top}\Xi^{\top}V_{xx}^{\prime}\Xi f_{x}\\ Q_{uu}&=\ell_{uu}+f_{u}^{\top}\Xi^{\top}V_{xx}^{\prime}\Xi f_{u}\\ Q_{xu}&=\ell_{xu}+f_{u}^{\top}\Xi^{\top}V_{xx}^{\prime}\Xi f_{x}\\ \end{split}$ (7)
@@ 4.1 DBaS-HiLQR
For smooth dynamical systems, the next barrier state $w_{k+1}$ is typically calculated by evaluating (12) at $x_{k+1}$$=$$f_{\Delta}(x_{k},u_{k},\Delta t)$ [11]. For hybrid dynamical systems, due to the necessity of assessing transitions via the saltation matrix, we modify the barrier state dynamics by using its continuous-time derivative $\dot{w}(t)$ as follows:
@@ 4.2 AL-HiLQR
    Fix $\lambda$ and $\mu$ parameters constant and solve the unconstrained problem for objective function (16) using HiLQR algorithm described in the beginning of Section 4 considering the modifications in the value function at the terminal state:
    $\begin{split}&V_{x}=(\ell_{N})_{x}+((g_{N})_{x})^{T}(\lambda+I_{\mu N}g_{N})\\ &V_{xx}=(\ell_{N})_{xx}+((g_{N})_{x})^{T}I_{\mu N}(g_{N})_{x}\\ \end{split}$ (18)
@@
    $\begin{split}&Q_{x}=\ell_{x}+f_{x}^{T}\Xi^{\top}V_{x}^{\prime}+g_{x}^{T}(\lambda+I_{\mu}g)\\ &Q_{u}=\ell_{u}+f_{u}^{T}\Xi^{\top}V_{x}^{\prime}+g_{u}^{T}(\lambda+I_{\mu}g)\\ &Q_{xx}=\ell_{xx}+f_{x}^{T}\Xi^{\top}V_{xx}^{\prime}\Xi f_{x}+g_{x}^{T}I_{\mu}g_{x}\\ &Q_{uu}=\ell_{uu}+f_{u}^{T}\Xi^{\top}V_{xx}^{\prime}\Xi f_{u}+g_{u}^{T}I_{\mu}g_{u}\\ &Q_{xu}=\ell_{xu}+f_{u}^{T}\Xi^{\top}V_{xx}^{\prime}\Xi f_{x}+g_{u}^{T}I_{\mu}g_{x}\\ \end{split}$ (19)
@@ 5 Experimental Results
In all test cases, no reference trajectory was provided. The initial and goal positions were $x_{0}$$=$$[0,4]^{\top}$ and $x_{goal}$$=$$[10,1]^{\top}$, respectively. The ball’s mass was $m$$=$$1$ $kg$, gravity was $g$$=$$9.81$ $m/s^{2}$, simulated with $N$$=$$200$ and $\Delta t$$=$$0.02$ $sec$.
For both methods the weight matrices have been taken equal to $Q_{T}=4\times 10^{3}I_{2\times 2}$ and $R=5\times 10^{-3}I_{2\times 2}$. For the velocity component, the weights have been set to 00. For the DBaS-HiLQR the barrier state weight has been chosen as $q_{w}=1\times 10^{-5}$. The solver archives convergence if all constraint tolerances are met and the error norm on the final position $\epsilon$ is $||\epsilon||\leq 5\times 10^{-2}$.
5.1 Test Case 1: Feasible starting rollout trajectory
@@
5.2 Test Case 2: Infeasible Starting Rollout Trajectory
@@
Table 1: Average Success Rates Across All Experiments
@@
Table 1 presents the average success rates across all experiments, considering all numbers of obstacles, and compares the performance of DBaS and AL in terms of feasible and infeasible starts. The results show that AL-HiLQR outperforms DBaS-HiLQR in handling infeasible starts, with only a $20.6\%$ drop in success rate, compared to $26.5\%$ for DBaS-HiLQR. AL-HiLQR performs better in handling infeasibilities due to its adaptability to additional constraints. In contrast, DBaS-HiLQR’s safety-embedded approach struggles when operating outside the safe set, leading to instability and a $23.4\%$ failure rate due to numerical issues.

### v2

@@ Abstract
I Introduction
@@ II Related Work
Direct methods transcribe the Optimal Control Problem (OCP) into a Nonlinear Programming (NLP) problem, and have been widely employed in robotic trajectory planning and control. However, hybrid dynamics involving continuous dynamics and discrete events often necessitate Linear Complementarity Programs (LCP) or Mixed Integer Programs (MIP) [9, 13]. Pardo et al. [14] presented an algorithm for legged robots’ locomotion, employing direct collocation within the constraint-consistent subspace defined by contact configurations. This method, incorporating projected impact dynamics constraints, was validated on a hydraulically-actuated quadruped robot. Posa et al. [6] introduced DIRCON, extending direct collocation with manifold constraints and third-order integration accuracy, demonstrating its efficacy in complex locomotion tasks. While these methods predefine a mode or gait sequence, limiting scalability, contact-implicit frameworks have improved optimization capabilities. Patel et al. [13] improved trajectory accuracy by combining direct collocation using higher-order orthogonal polynomials with contact-implicit optimization, eliminating the need for mode scheduling and a priori contact knowledge. Posa et al. [9] addressed trajectory planning with inelastic contact models formulated as LCP, solved via Sequential Quadratic Programming, achieving accurate contact-implicit trajectory optimization. Earlier, Yunt [15] proposed an Augmented Lagrangian-based direct shooting method for switching Lagrangian systems, eliminating the need for explicit mode scheduling by formulating the NLP in the mathematical programming with equilibrium constraints framework. Additionally, Manchester et al. [16] developed algorithms based on discrete variational mechanics for higher-order time-stepping methods, enhancing the accuracy of contact-implicit trajectory optimization.
@@ IV Hybrid iLQR
[…]
@@ IV-A DBaS-HiLQR
For smooth dynamical systems, the next barrier state $w_{k+1}$ is typically calculated by evaluating (12) at $x_{k+1}$$f_{\Delta}(x_{k},u_{k},\Delta t)$ [11]. For hybrid dynamical systems, due to the necessity of assessing transitions via the saltation matrix, we modify the barrier state dynamics by using its continuous-time derivative $\dot{w}(t)$ as follows:
@@ IV-B AL-HiLQR
    Fix $\lambda$ and $\mu$ parameters constant and solve the unconstrained problem for objective function (16) using HiLQR algorithm described in the beginning of Section IV considering the modifications in the value function at the terminal state:
[…]
@@
[…]
@@ V Experimental Results
In all test cases, no reference trajectory was provided. The initial and goal positions were $x_{0}$$0,4]^{\top}$ and $x_{goal}$$10,1]^{\top}$, respectively. The ball’s mass was mm==11 kgkg, gravity was gg==$9.81$ $m/s^{2}$, simulated with NN==$200$ and $\Delta t$==$0.02$ secsec.
For both methods the weight matrices have been taken equal to $Q_{T}=4\times 10^{3}I_{2\times 2}$ and $R=5\times 10^{-3}I_{2\times 2}$. For the velocity component, the weights have been set to 0. For the DBaS-HiLQR the barrier state weight has been chosen as $q_{w}=1\times 10^{-5}$. The solver archives convergence if all constraint tolerances are met and the error norm on the final position $\epsilon$ is $\epsilon||\leq 5\times 10^{-2}$.
V-A Test Case 1: Feasible starting rollout trajectory
@@
V-B Test Case 2: Infeasible Starting Rollout Trajectory
@@
TABLE I: Average Success Rates Across All Experiments
@@
Table I presents the average success rates across all experiments, considering all numbers of obstacles, and compares the performance of DBaS and AL in terms of feasible and infeasible starts. The results show that AL-HiLQR outperforms DBaS-HiLQR in handling infeasible starts, with only a $20.6\%$ drop in success rate, compared to $26.5\%$ for DBaS-HiLQR. AL-HiLQR performs better in handling infeasibilities due to its adaptability to additional constraints. In contrast, DBaS-HiLQR’s safety-embedded approach struggles when operating outside the safe set, leading to instability and a $23.4\%$ failure rate due to numerical issues.